OpenStudy (anonymous):

solving trig equations: 1. cos θ = 1/4 2. sin θ = -0.4 3. sin θ = -0.15 4. tan θ = -1

OpenStudy (anonymous):

Do you want to show the exact solution or will a decimal approximation from a calculator suffice?

OpenStudy (anonymous):

For all of these, the unknown to solve for is Θ. It is found by taking the inverse trig function of the given ratio.

OpenStudy (anonymous):

exact solution is best. I get that its the inverse but I have to show it in general form like this: \[\pi/2 +2k \pi \]

OpenStudy (anonymous):

I see. Ok, that's fine. #4. is the easiest to see right away. The tangent of an angle is equal to the slope of the line. What angle is formed by a line with a slope of -1?

OpenStudy (anonymous):

\[\theta =\cos^{-1} (1/4)\]

OpenStudy (anonymous):

\[\theta=\sin^{-1} (-0.4)\]

OpenStudy (anonymous):

\[\theta=\sin^{-1} (-0.15)\]

OpenStudy (anonymous):

\[\theta= \tan^{-1} (-1)\]

OpenStudy (anonymous):

solve them by your calculator to find answers... if not then tell me i'll send you the answers too..

OpenStudy (anonymous):

I understand how to do that, but I don't get how to have them in \[\pi \] form?

OpenStudy (anonymous):

multiply answers by (pi/180) you'll get what you need... then answers will be in radians..

OpenStudy (anonymous):

look angle in "pi" means angle in radian not in degree... so you can convert the degree angles into radians by multiplying it by (pi/180).. got it now??

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

Only problem with that is if you use a calculator to evaluate, say, arccos(1/4) = 1.318116... radians. It is difficult to see how many multiples or parts of π that is.

OpenStudy (anonymous):

I get what your saying and thats why its not accepting the answers for my homework